OpenStudy (anonymous):

solve:2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

open up the parenthesis: \[2x-6 \ge 5-x-3\] can you proceed from here?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

OK, first get the like terms to one side, we can put the unknowns(those with x) on the LHS and the knowns on the RHS, such that: \[2x+x \ge 5-3+6\] got it up to there?

OpenStudy (anonymous):

I think I did it wrong, I got 2.67

OpenStudy (anonymous):

you're somehow right, but this is an inequality, so the answer should be: \[x \ge 2.67\]

OpenStudy (anonymous):

Cool! Thanks.

OpenStudy (anonymous):

\[3x \ge 8\] divide all through by 3 \[x \ge 2.67\]